# PoseGravity: Pose Estimation from Points and Lines with Axis Prior

### v1

@@ 1 Introduction
Estimating camera pose from image features and known 3D geometry is crucial for many computer vision applications such as augmented reality, visual odometry, and structure from motion. In certain scenarios, prior knowledge of the camera’s rotation axis is available (e.g. IMU measurement, detected vanishing point, or domain assumptions). Using this prior knowledge, the resulting problem reduces from six degrees of freedom (three for orientation, three for position) to four degrees of freedom as the rotation matrix only has only degree of freedom to rotate about the given axis. Some benefits of this simplified problem include reducing the minimum correspondences needed, improving the robustness of estimation to noise, and providing a quick initial solution for the general estimation problem. Previous solutions have been proposed that solve this problem for specific types of configurations or correspondences. In this paper, a fast and accurate algorithm is presented that generalizes the solution further by solving the problem for both point and line feature correspondences in minimal and overconstrained cases.
@@ 3.2 Constrained Least Squares Formulation
$\mathbf{T}=\mathbf{S}\mathbf{r}$
$=(\sum_{i}\mathbf{Q}_{i}^{p}+\sum_{j}\mathbf{Q}_{j}^{l})^{-1}(\sum_{i}\mathbf{Q}_{i}^{p}\mathbf{A}_{i}^{p}+\sum_{j}\mathbf{Q}_{j}^{l}\mathbf{A}_{j}^{l})\mathbf{r}$ (17)
@@ 4.1 Experiment Setup
To simulate noisy measurements and detections, gravity vector measurement noise and 2D detection noise are added. For the former, the gravity vector measurement is rotated around a random axis by a normally distributed angle $\theta_{noise}$ in degrees. For the latter, independent Gaussian noise $\epsilon_{noise}$ is added to each component of each 2D feature.
@@ A.2 Loss Conic Rank for Minimal Problems
Proposition 2: For any nonzero, linearly independent, and minimal ($n+m=2$, $n\geq 1$) inputs, $rank(\mathbf{\Omega})\leq 1$.
Proof It can be assumed without loss of generality that image point $\mathbf{p}_{0}$ corresponds to the 3D point $\mathbf{d}_{0}=(0,0,0)$ since the translation estimate is relative and each configuration has at least one point that can define the new coordinate frame origin.
Consider the problem with $n=2$ points $\mathbf{p}_{0}$ and $\mathbf{p}_{1}$. The expression for loss conic can be expanded and simplified to the following:
$\mathbf{\Omega}=(\mathbf{A}_{1}^{p})^{T}\mathbf{Q}_{1}^{p}\mathbf{A}_{1}^{p}-(\mathbf{Q}_{1}^{p}\mathbf{A}_{1}^{p})^{T}(\mathbf{Q}_{0}^{p}+\mathbf{Q}_{1}^{p})^{-1}\mathbf{Q}_{1}^{p}\mathbf{A}_{1}^{p}$
$=(\mathbf{A}_{1}^{p})^{T}(\mathbf{Q}_{1}^{p}-\mathbf{Q}_{1}^{p}(\mathbf{Q}_{0}^{p}+\mathbf{Q}_{1}^{p})^{-1}\mathbf{Q}_{1}^{p})\mathbf{A}_{1}^{p}$
$=(\mathbf{A}_{1}^{p})^{T}\mathbf{K}\mathbf{A}_{1}^{p}$
Examining the matrix $\mathbf{K}$, we see that $\mathbf{p}_{1}\in ker(\mathbf{K})$. Nontrivially, we can verify by calculation that $\mathbf{p}_{0}\in ker(\mathbf{K})$ as well. Since the two points are assumed linearly independent, $dim(ker(\mathbf{K}))\geq 2$, so $rank(\mathbf{K})\leq 1$. Multiplication by $\mathbf{A}_{1}^{p}$ can only maintain or decrease the rank of the product, so $rank((\mathbf{A}_{1}^{p})^{T}\mathbf{K}\mathbf{A}_{1}^{p})=rank(\mathbf{\Omega})\leq 1$.
Consider the problem with $n=1$ point $\mathbf{p}_{0}$ and $m=1$ line $\mathbf{n}_{0}$. The expression for loss conic similarly can be expanded and simplified to the following:
$\mathbf{\Omega}=(\mathbf{A}_{0}^{l})^{T}(\mathbf{Q}_{0}^{l}-\mathbf{Q}_{0}^{l}(\mathbf{Q}_{0}^{p}+\mathbf{Q}_{0}^{l})^{-1}\mathbf{Q}_{0}^{l})\mathbf{A}_{0}^{l}+\mathbf{V}_{0}^{T}\mathbf{Q}_{0}^{l}\mathbf{V}_{0}$
$=(\mathbf{A}_{0}^{l})^{T}\mathbf{L}\mathbf{A}_{0}^{l}+\mathbf{V}_{0}^{T}\mathbf{Q}_{0}^{l}\mathbf{V}_{0}$
Nontrivially, we can verify by calculation that $\mathbf{L}$ is the zero matrix. $rank(\mathbf{Q}_{0}^{l})=1$ and multiplication by $\mathbf{V}_{0}$ can only maintain or decrease rank, so $rank(\mathbf{V}_{0}^{T}\mathbf{Q}_{0}^{l}\mathbf{V}_{0})=rank(\mathbf{\Omega})\leq 1$.
@@
Intuitively, this result stems from the fact that minimal configurations constrain four degrees of freedom while the estimation problem has six degrees of freedom. Estimation of $\mathbf{T}$ consumes three of those constraints, so there is only one left when constructing $\mathbf{\Omega}$. The problem is solved by applying the additional constraints to solution $\mathbf{r}$ that $x^{2}+y^{2}=1$ and the third element of $\mathbf{r}$ is 1. Also note that the eigenvector corresponding to the nonzero eigenvalue of $\mathbf{K}$ when $rank(\mathbf{K})=1$ is $\mathbf{p}_{0}\times\mathbf{p}_{1}$ which defines the projective line between the points, playing a similar role to $\mathbf{n}_{0}$ in $\mathbf{Q}_{0}^{l}$.
@@ A.3 Recovery of Non-Exact Minimal Solutions
Proposition 3: If the zero level curve of a rank one loss conic in the x-y plane exists and has no real solutions with the unit circle, then the best feasible loss-minimizing solution is the point on the circle closest to the curve.
@@
The distance squared from a point to a line is the original loss function scaled by a constant. Therefore, minimizing the distance (and thereby distance squared) minimizes the loss function. Since the line is assumed to not intersect any point on the unit circle, the best feasible solution is the point on the circle closest to the line. Obtaining $l$ from conic decomposition, this solution can be simply calculated as:
@@
If $a=b=0$, the conic does not exist in the $x{\text{-}}y$ plane and a uniquely loss minimizing solution cannot be found.
@@ Appendix B Note on Line Loss Scale $\delta$
For line features, the proposed solver’s loss function combines the rotation and translation loss terms instead of estimating them sequentially like other approaches. However, since those two terms in loss function in equation 16 have different 3D objects associated with them, there can be a bias from the points $\mathbf{m}_{j}$ and the line directions $\mathbf{v}_{j}$ having different scales (e.g. if $\mathbf{v}_{j}$ is normalized), leading to the loss being dominated by one term and resulting in suboptimal pose estimates. Estimates from planar configurations are particularly sensitive to this scale difference. This issue does not occur when dealing only with point features as the loss is encapsulated in a single term that includes both rotation and translation components in a naturally balanced way. To compensate for the bias, we found it sufficient to introduce a constant scaling factor $\delta^{2}$ which manually balances the line direction loss. In our experiments, we use $\delta=100$ with normalized directions $\mathbf{v}_{j}$ for standardization.

### v2

@@ 1 Introduction
Estimating camera pose from image features and known 3D geometry is crucial for many computer vision applications such as augmented reality, visual odometry, and structure from motion. In certain scenarios, prior knowledge of the camera’s rotation axis is available (e.g. IMU measurement, detected vanishing point, or domain assumptions). Using this prior knowledge, the resulting problem reduces from six degrees of freedom (three for orientation, three for position) to four degrees of freedom as the rotation matrix only has one degree of freedom to rotate about the given axis. Some benefits of this simplified problem include reducing the minimum correspondences needed, improving the robustness of estimation to noise, and providing a quick initial solution for the general estimation problem. Previous solutions have been proposed that solve this problem for specific types of configurations or correspondences. In this paper, a fast and accurate algorithm is presented that generalizes the solution further by solving the problem for both point and line feature correspondences in minimal and overconstrained cases.
@@ 3.2 Constrained Least Squares Formulation
$\mathbf{T}=\mathbf{S}\mathbf{r}=$
$-(\sum_{i}\mathbf{Q}_{i}^{p}+\sum_{j}\mathbf{Q}_{j}^{l})^{-1}(\sum_{i}\mathbf{Q}_{i}^{p}\mathbf{A}_{i}^{p}+\sum_{j}\mathbf{Q}_{j}^{l}\mathbf{A}_{j}^{l})\mathbf{r}$ (17)
@@ 4.1 Experiment Setup
To simulate noisy measurements and detections, gravity vector measurement noise and detection noise are added. For the former, the gravity vector measurement is rotated around a random axis by a normally distributed angle $\theta_{noise}$ in degrees. For the latter, independent Gaussian noise $\epsilon_{noise}$ is added to each component of each 2D feature.
@@ A.2 Loss Conic Rank for Minimal Problems
Proposition 2: For nonzero, linearly independent, and minimal ($n+m=2$, $n\geq 1$) inputs, $rank(\mathbf{\Omega})\leq 1$.
Proof For the case of two points, denoting $\mathbf{Q}_{inv}=(\mathbf{Q}_{0}^{p}+\mathbf{Q}_{1}^{p})^{-1}$, $\mathbf{\Omega}$ is expanded to the following terms:
$\mathbf{\Omega}=\mathbf{A}_{0}^{pT}\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p}+\mathbf{A}_{1}^{pT}\mathbf{Q}_{1}^{p}\mathbf{A}_{1}^{p}-$
$(\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p})^{T}\mathbf{Q}_{inv}\mathbf{Q}_{1}^{p}\mathbf{A}_{1}^{p}-(\mathbf{Q}_{1}^{p}\mathbf{A}_{1}^{p})^{T}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p}-$
$(\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p})^{T}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p}-(\mathbf{Q}_{1}^{p}\mathbf{A}_{1}^{p})^{T}\mathbf{Q}_{inv}\mathbf{Q}_{1}^{p}\mathbf{A}_{1}^{p}$
$=\mathbf{A}_{0}^{pT}(\mathbf{Q}_{0}^{p}-\mathbf{Q}_{0}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p})\mathbf{A}_{0}^{p}-$
$\mathbf{A}_{1}^{pT}\mathbf{Q}_{1}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p}-\mathbf{A}_{0}^{pT}\mathbf{Q}_{0}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{1}^{p}\mathbf{A}_{1}^{p}+$
$\mathbf{A}_{1}^{pT}(\mathbf{Q}_{1}^{p}-\mathbf{Q}_{1}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{1}^{p})\mathbf{A}_{1}^{p}$
$=\mathbf{A}_{0}^{pT}\mathbf{K}_{0}\mathbf{A}_{0}^{p}-\mathbf{A}_{1}^{pT}\mathbf{K}_{1}\mathbf{A}_{0}^{p}-\mathbf{A}_{0}^{pT}\mathbf{K}_{2}\mathbf{A}_{1}^{p}+\mathbf{A}_{1}^{pT}\mathbf{K}_{3}\mathbf{A}_{1}^{p}$
We next examine the inner $\mathbf{K}_{i}$ matrices. Taking the difference of the first one with the others yields:
$\mathbf{K}_{0}-\mathbf{K}_{1}$ $=(\mathbf{Q}_{0}^{p}-\mathbf{Q}_{0}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p})-\mathbf{Q}_{1}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p}$
$=\mathbf{Q}_{0}^{p}-(\mathbf{Q}_{0}^{p}+\mathbf{Q}_{1}^{p})\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p}=\mathbf{Q}_{0}^{p}-\mathbf{Q}_{0}^{p}=0,$
$\mathbf{K}_{0}-\mathbf{K}_{2}$ $=(\mathbf{Q}_{0}^{p}-\mathbf{Q}_{0}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p})-\mathbf{Q}_{0}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{1}^{p}$
$=\mathbf{Q}_{0}^{p}-\mathbf{Q}_{0}^{p}\mathbf{Q}_{inv}(\mathbf{Q}_{0}^{p}+\mathbf{Q}_{1}^{p})=\mathbf{Q}_{0}^{p}-\mathbf{Q}_{0}^{p}=0,$
$\mathbf{K}_{0}-\mathbf{K}_{3}$ $=\mathbf{K}_{0}-(\mathbf{Q}_{1}^{p}-\mathbf{Q}_{1}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{1}^{p})$
$=\mathbf{K}_{0}-\mathbf{Q}_{1}^{p}+\mathbf{Q}_{1}^{p}\mathbf{Q}_{inv}(\mathbf{Q}_{0}^{p}+\mathbf{Q}_{1}^{p})-\mathbf{K}_{1}$
$=\mathbf{K}_{0}-\mathbf{Q}_{1}^{p}+\mathbf{Q}_{1}^{p}-\mathbf{K}_{1}=0$
Thus, all the inner matrices are equivalent ($\mathbf{K}_{i}=\mathbf{K}$). Furthermore, we can examine $\mathbf{K}_{1}$ or $\mathbf{K}_{2}$ to see that the two image points $\mathbf{p}_{0},\mathbf{p}_{1}\in ker(\mathbf{K})$ trivially. Since the points are assumed linearly independent, $dim(ker(\mathbf{K}))\geq 2$, so the rank of $\mathbf{K}$ is at most 1. We can now rewrite the loss as:
$\mathbf{\Omega}=(\mathbf{A}_{0}^{p}-\mathbf{A}_{1}^{p})^{T}\mathbf{K}(\mathbf{A}_{0}^{p}-\mathbf{A}_{1}^{p})$
Since $\mathbf{K}$ has at most rank 1, and multiplying by other matrices cannot increase the rank, $rank(\mathbf{\Omega})\leq 1$.
For the case of one point and one line, denoting $\mathbf{Q}_{inv}=(\mathbf{Q}_{0}^{p}+\mathbf{Q}_{0}^{l})^{-1}$, the terms of $\mathbf{\Omega}$ are:
$\mathbf{\Omega}=\mathbf{A}_{0}^{pT}\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p}+\mathbf{A}_{0}^{lT}\mathbf{Q}_{0}^{l}\mathbf{A}_{0}^{l}+\delta^{2}\mathbf{V}_{0}^{T}\mathbf{Q}_{0}^{l}\mathbf{V}_{0}-$
$(\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p})^{T}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p}-(\mathbf{Q}_{0}^{l}\mathbf{A}_{0}^{l})^{T}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p}-$
$(\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p})^{T}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{l}\mathbf{A}_{0}^{l}-(\mathbf{Q}_{0}^{l}\mathbf{A}_{0}^{l})^{T}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{l}\mathbf{A}_{0}^{l}$
$=\mathbf{A}_{0}^{pT}(\mathbf{Q}_{0}^{p}-\mathbf{Q}_{0}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p})\mathbf{A}_{0}^{p}-$
$\mathbf{A}_{0}^{lT}\mathbf{Q}_{0}^{l}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{p}\mathbf{A}_{0}^{p}-\mathbf{A}_{0}^{pT}\mathbf{Q}_{0}^{p}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{l}\mathbf{A}_{0}^{l}+$
$\mathbf{A}_{0}^{lT}(\mathbf{Q}_{0}^{l}-\mathbf{Q}_{0}^{l}\mathbf{Q}_{inv}\mathbf{Q}_{0}^{l})\mathbf{A}_{0}^{l}+\delta^{2}\mathbf{V}_{0}^{T}\mathbf{Q}_{0}^{l}\mathbf{V}_{0}$
$=\mathbf{A}_{0}^{pT}\mathbf{L}_{0}\mathbf{A}_{0}^{p}-\mathbf{A}_{0}^{lT}\mathbf{L}_{1}\mathbf{A}_{0}^{p}-\mathbf{A}_{0}^{pT}\mathbf{L}_{2}\mathbf{A}_{0}^{l}+$
$\mathbf{A}_{0}^{lT}\mathbf{L}_{3}\mathbf{A}_{0}^{l}+\delta^{2}\mathbf{V}_{0}^{T}\mathbf{Q}_{0}^{l}\mathbf{V}_{0}$
Performing the same differences between $\mathbf{L}_{0}$ and the other $\mathbf{L}_{i}$ matrices as in the two point case, we arrive at the similar result that all the $\mathbf{L}_{i}$ matrices are equivalent ($\mathbf{L}_{i}=\mathbf{L}$). Examining $\mathbf{L}_{1}$ or $\mathbf{L}_{2}$, we see that the image point and the image line’s orthogonal complement $\mathbf{p}_{0},\mathbf{n}_{0}^{\bot}\in ker(\mathbf{L})$ trivially and are one and two dimensional respectively. Since the image objects are assumed linearly independent (i.e. $\mathbf{p}_{0}\cdot\mathbf{n}_{0}\neq 0$), the point together with the line’s orthogonal complement form a three dimensional basis for the nullspace of $\mathbf{L}$ implying that $\mathbf{L}$ is the zero matrix. As a result, the only remaining term in the expression is the direction loss $\delta^{2}\mathbf{V}_{0}^{T}\mathbf{Q}_{0}^{l}\mathbf{V}_{0}$. Since $\mathbf{Q}_{0}^{l}$ has rank 1, and multiplying by other matrices and scaling cannot increase the rank, $rank(\mathbf{\Omega})\leq 1$.
@@
Intuitively, this result stems from the fact that minimal configurations constrain four degrees of freedom while the estimation problem has six degrees of freedom. Estimation of $\mathbf{T}$ consumes three of those constraints, so there is only one left when constructing $\mathbf{\Omega}$. The problem is solved by applying additional constraints to the solution $\mathbf{r}$ specifying that $x^{2}+y^{2}=1$ and the third element of $\mathbf{r}$ is 1. It is noteworthy that in the two-point case where $rank(\mathbf{K})=1$, the eigenvector associated with the nonzero eigenvalue of $\mathbf{K}$ is given by $\mathbf{p}_{0}\times\mathbf{p}_{1}$. This vector defines the projective line passing through the two points analogous to $\mathbf{n}_{0}$. In addition, $\mathbf{A}_{0}^{p}-\mathbf{A}_{1}^{p}$ defines the direction of the line passing through the 3D points analogous to $\mathbf{V}_{0}$. Thus, $\mathbf{\Omega}$ has the same geometric interpretation in either case.
@@ A.3 Recovery of Non-Exact Minimal Solutions
Proposition 3: If the zero level curve of a rank 1 loss conic in the x-y plane exists and has no real solutions with the unit circle, then the best feasible solution is the point on the circle closest to the curve.
@@
The distance squared from a point to a line is the original loss function scaled by a constant. Therefore, minimizing the distance (and thereby distance squared) minimizes the loss function. As the line is assumed to not intersect the unit circle, the point on the circle closest to the line is the loss-minimizing solution. Obtaining $l$ from conic decomposition, this solution can be simply calculated as:
@@
If $a=b=0$, the conic does not exist in the $x{\text{-}}y$ plane, and a unique loss-minimizing solution cannot be found.
@@ Appendix B Note on Line Loss Scale $\delta$
For line features, the proposed solver’s loss function combines the rotation and translation loss terms instead of estimating them sequentially like other approaches. Since those two terms in loss function in equation 16 have different 3D objects associated with them, there can be a bias from the points $\mathbf{m}_{j}$ and the line directions $\mathbf{v}_{j}$ having different scales (e.g. if $\mathbf{v}_{j}$ is normalized), leading to the loss being dominated by one term and resulting in suboptimal pose estimates. Estimates from planar configurations are particularly sensitive to this scale difference. This issue does not occur when dealing only with point features as the loss is encapsulated by a single term which includes both rotation and translation components in a naturally balanced way. To compensate for the bias, we found it sufficient to introduce a constant scaling factor $\delta^{2}$ which manually balances the line direction loss. In our experiments, we use $\delta=100$ with normalized directions $\mathbf{v}_{j}$ for standardization.